# SKYSURF IX - The Cosmic Optical and Infrared Background from Integrated Galaxy Light Measurements

### v1

@@ Abstract
As part of the SKYSURF Hubble Space Telescope (HST) Legacy Archival program we present galaxy number counts which yield measurements of the extragalactic background light (EBL) at 15 different wavelengths. We have processed 82,752 HST images across 23 filters into 16,686 mosaics using the same software and processing pipeline throughout. Using 17/23 filters that give reliable galaxy counts, we constrain the integrated galaxy light (IGL) with a 1.5-9% error between 0.3 and 1.6 m in combination with 8 bands from WAVES (Wide Area VISTA Extragalactic Survey) and DEVILS (Deep Extragalactic Visible Legacy Survey). While HST was never intended to undertake large area surveys, through extensive quality control and filtering, we were able to extract a reliable and representative sample of fields distributed across the sky. Our final catalogs cover a combined , with individual filters covering areas ranging from . The combination of numerous independent sight-lines and area coverage allows us to reduce cosmic variance uncertainties in deep number counts to 0.06%-1.8%. For the first time we are able to establish a measurement of the IGL, , at 0.59 m using HST data. We obtain a cosmic optical background value of . Different techniques used to measure the COB, both directly and indirectly, have recently converged indicating that the COB arises almost exclusively from processes within galaxies. This in combination with the recent values reported from New Horizons and very high energy (VHE) constraints leaves very little room for any diffuse emission coming from outside the Milky Way.
1 Introduction
@@
The extragalactic background light (EBL) is composed of all radiation originating from outside of our galaxy (Hill et al., 2018; Driver, 2021). The EBL is typically broken down into wavelength regimes such as the cosmic optical background (COB ) , cosmic infrared background (CIB ) and cosmic radio background (CRB); e.g. (Hill et al., 2018). The cosmic microwave background (CMB) is distinct from other astrophysical backgrounds as it is a relic from when the early Universe cooled to a temperature where neutral atoms could form. The remaining EBL from -ray to radio wavelengths contains information about all physical processes that have produced photons since recombination and thus can be considered separate from the CMB. The COB and CIB dominate the EBL and provide roughly equal energy contributions (Driver et al., 2016b; Koushan et al., 2021), though they both respectively fall a factor of 40x below the intensity of the CMB. The COB is perhaps the most important component of the EBL due to its intimate relationship with the cosmic star formation history (CSFH). The CSFH is defined as the volume averaged rate at which stars form as a function of either look-back time or redshift, for a through review see (Madau & Dickinson, 2014). The COB can be directly predicted from the CSFH assuming an implementation of a stellar initial mass function (IMF), dust attenuation model, metallicity evolution history, and stellar population model which has been demonstrated in works such as Andrews et al. (2017) and Koushan et al. (2021). Thus, empirical constraints of the COB can be used to test these models. The contribution from AGN is also important; especially at UV and mid-IR wavelengths, and is also able to be modeled. The cosmic AGN luminosity history (CAGNH) is known to co-evolve with the CSFH (D’Silva et al., 2023). Despite its importance, the COB remains one of the most difficult portions of the EBL to study in large part due to the presence of foregrounds which outshine the EBL, see (Bernstein et al., 2002; Windhorst et al., 2022; Caddy et al., 2022; Carleton et al., 2022; Windhorst et al., 2023).
@@
Indirect detection techniques have become possible with advancements in the ability to study high energy -ray sources. The interaction between photons from distant -ray sources, typically blazars, and optical EBL photons results in an attenuation of TeV energy -rays through particle/anti-particle pair production interactions as discussed in Desai et al. (2017). Using blazar spectra with established redshifts, they are able to estimate the attenuation of -ray spectra by EBL photons along the line of sight between the source and detectors. While difficult, this method (hereafter VHE) is less susceptible to foregrounds and is also unaffected by issues such as cosmic variance and zero point uncertainties, which plague the IGL. However VHE techniques typically apply a universal spectral shape to the -ray sources and infer deviations from this spectrum as entirely due to interactions with EBL photons or intergalactic magnetic fields. The technique typically requires an existing COB model and just constraints the normalization of thhe model. Recent developments in VHE studies of the EBL, (e.g. Gréaux & Biteau, 2023; Gréaux et al., 2024) have adopted a Bayesian framework to avoid limitations introduced by assuming a model and spectral shape for both the EBL and gamma ray source. Using the observed -ray spectra of blazars they present preliminary results in which they now extract the COB shape and amplitude. Uncertainties in measurements of the COB established by VHE studies to 15% and will be developed further in the coming years.
@@
A commonly employed method to place constraints on the EBL without accounting for foregrounds is to use galaxy number counts and measure the total intensity of all resolved galaxies in works such as (e.g. Madau & Pozzetti, 2000; Totani et al., 2001; Xu et al., 2005; Dole et al., 2006; Hopwood et al., 2010; Keenan et al., 2010; Berta et al., 2011; Béthermin et al., 2012; Driver et al., 2016b; Koushan et al., 2021; Windhorst et al., 2022, 2023, Carter et al. 2025(submitted)). The IGL can be well constrained by surveying a large enough volume to obtain a representative sample of the Universe. Galaxy number counts, or rather the galaxy flux density, can be integrated to provide a measurement of their combined intensity. While establishing robust lower limits, the IGL is prone to uncertainties introduced by cosmic variance and magnitude zero-point differences present between the facilities providing the number counts. Cosmic variance and zero point errors remain the dominant sources of uncertainty in measuring the IGL. In work such as Driver et al. (2016b) these contribute uncertainties of 10% and 5%, respectively. More recent studies such as Koushan et al. (2021), through the use of larger area surveys and updated zero point calibrations, were able to bring these uncertainties down to 4%. However, Driver et al. (2016b) and Koushan et al. (2021) relied only on a small number of deep HST fields. Work such as Windhorst et al. (2011) and Rafelski et al. (2015) provide the faintest optical galaxy number counts and have only been surpassed by JWST recently in the NIR (Windhorst et al., 2023). The two HST surveys included in Driver et al. (2016b) and Koushan et al. (2021) cover a combined 40-60 arcmin2, and thus cosmic variance remains the dominant source of uncertainty in these surveys at 10-15% as estimated using the formula developed in Driver & Robotham (2010). Uncertainties in the IGL have also been reduced to between by Carter et al. 2025 (submitted). SKYSURF aims to use a significant fraction of the area covered by HST over 20 years of observations to further reduce uncertainties in measurements of the IGL primarily by reducing the cosmic variance errors of individual HST deep surveys used in previous works.
@@
Despite efforts to measure the COB both directly and through the IGL, disagreements between the upper limits established by work of Carleton et al. (2022) O’Brien et al. (2023), Mattila et al. (2017), Lauer et al. (2022) and lower limits established by the IGL in Driver et al. (2016b) and Koushan et al. (2021) remained factors of apart. The newest direct measurements of Postman et al. (2024) from New Horizons suggest that the tension between the IGL and direct measurements has finally been resolved. Again using the LORRI instrument aboard New Horizons, measurements have yielded a significant detection of the EBL at the LORRI pivot wavelength of , with the main change in Postman et al. (2024) being an improved method for estimating the DGL. Within their respective errors, the measurements of Postman et al. (2024), Koushan et al. (2021), and preliminary VHE measurements from Gréaux & Biteau (2023) have finally converged. The explanation posed in Postman et al. (2024) is that the IGL composes the most significant component of the COB and the majority of extragalactic optical flux originates from within galaxies. The JWST work of Windhorst et al. (2023) suggests that no more than 20% of the COB can originate from outside of galaxies. As such, efforts to reduce uncertainties in the IGL can better constrain models and provide a near complete census of photon production from baryon-photon decoupling until the present day.
@@ 2.1 The WAVES Survey
WAVES is a spectroscopic redshift survey to be conducted using the European Southern Observatory’s 4-Metre-Multi-Object Spectroscopic Telescope (4MOST) (de Jong, 2011). In this work, bright galaxy number counts are produced in preparation for the WAVES survey (Driver et al., 2016a, 2019, Bellstedt et al. in prep) which covers a region of the sky. The galaxy catalogs used for WAVES target selection (Bellstedt et al. in prep) are produced from imaging taken from the VISTA KIlo-degree Infrared Galaxy (Edge et al., 2013, VIKING) and the VST Kilo-Degree Survey (Kuijken et al., 2019, VST-KiDS), both of which are ESO public data sets. The data was WCS aligned using ProPane (Robotham et al., 2023), and object detection is done using an rizY inverse variance stack using ProFound (Robotham et al., 2018). All bright galaxies were inspected for fragmentation following the detection phase and fixed. Following the detection phase, fluxes in each band were determined using forced aperture photometry from the detection stack. Stars brighter than GAIA g mag were masked. Star-galaxy separation was achieved using an unsupervised machine learning technique developed in Cook et al. (2024). See Bellstedt et al. (in prep) for full details on these steps. In this work we use WAVES input catalog v1.2 with 7 bands (ugrizYJH) that overlap with the HST filter system. The completeness limit of each band is shown in Table 1 where completeness limits were defined by stepping back by one half-magnitude bin from where the counts reach a maximum.
@@ 2.2 The DEVILS Survey
For intermediate magnitude number counts we turn to DEVILS (Davies et al., 2021), using all three DEVILS fields, D02, D03, and D10. The three fields are centered on well-studied regions, XMM-LSS, ECDFS, and COSMOS, respectively. The DEVILS galaxy number counts are complete to AB 24 mag across all 3 fields and cover a combined area of 6 . A brief summary of areas and completeness limits are provided in Table 1. We elect to use the same 7 bands chosen from the WAVES survey, i.e. , . We note that D02 and D10 were observed by the Hyper Suprime Cam (HSC) (Aihara et al., 2017) in griz with the YJH bands observed by VISTA. In D03 the griz imaging was obtained by the VST. Finally, u-band imaging was obtained from the Canada-France-Hawaii-Telescope (CFHT) and VST.
@@ 2.3 The SKYSURF HST Archival Program
SKYSURF is an HST cycle 27-29 Legacy Archival program (AR-09955 & AR-15810) designed to use a significant amount of HST imaging data to measure the sky-surface brightness (sky-SB) across the wavelength range covered by HST, i.e., . This requires the reprocessing of archived HST data using consistent software pipelines throughout. During the last two years, the SKYSURF project has measured the sky-SB and divided it into its components, the ZL, DGL, and EBL, (e.g. O’Brien et al., 2023; Carleton et al., 2022; Windhorst et al., 2022, Carter et al. 2025 (Submitted)). The SKYSURF project is split into two separate teams using independent processing pipelines and methodologies to obtain mosaics, sky-SB values, and object catalogs. This paper constitutes the portion of the project led from the University of Western Australia focused on integrated galaxy light (IGL) measurements using object catalogs generated using the ProPane and ProFound software packages, (Robotham et al., 2023, 2018) as was used by WAVES and DEVILS. In Carter et al. 2025 (submitted), an independent processing of the SKYSURF archive was performed using the AstroDrizzle (Gonzaga et al., 2012) for image alignment and stacking. Source detection was done using SourceExtractor (Bertin & Arnouts, 1996). Our overriding objective is to produce reliable galaxy number counts across as broad a magnitude and wavelength range as possible. We begin our SKYSURF selection by only including HST imaging with an exposure time of 200 seconds or longer and at galactic latitude . A projection showing the location of SKYSURF data used in this work is shown in Figure 1. All data was downloaded from the Mikulski Archive for Space Telescopes (MAST)111https://mast.stsci.edu/portal/Mashup/Clients/Mast/Portal.html.
@@ 3.1 Galactic Extinction Corrections
We correct for galactic extinction using the Planck satellite dust maps accessible through the web tool333https://irsa.ipac.caltech.edu/applications/DUST/ with extinction estimates provided by Finkbeiner et al. (1998). Given the small angular size of SKYSURF fields compared to the resolution of the dust maps, we elect to use the center of the field to obtain a single E(B-V) value and thus apply one extinction value to all objects in that field. Using the Spanish Virtual Observatory tool provided by (Rodrigo & Solano, 2020)444http://svo2.cab.inta-csic.es/theory/fps/, we obtain values for each HST filter. The Av values adopted for each HST filter used in this work along with the pivot wavelengths adopted in this work are listed in Table 3. In Figure 6 we show the Galactic extinction values for each mosaic as a function of Galactic latitude. By design, SKYSURF omitted all observations within 10 degrees of the galactic plane, as indicated by the vertical lines in Figure 6.

### v2

# SKYSURF IX - The Cosmic Optical Background from Integrated Galaxy Light Measurements

@@ Abstract
As part of the SKYSURF Hubble Space Telescope (HST) Legacy Archival program we present galaxy number counts which yield measurements of the extragalactic background light (EBL) at 15 different wavelengths. We have processed 82,752 HST images across 23 filters into 16,686 mosaics using the same software and processing pipeline throughout. Using 17/23 filters that give reliable galaxy counts, we constrain the integrated galaxy light (IGL) with a 1.1-9% error between 0.3 and 1.6 m in combination with 8 bands from WAVES (Wide Area VISTA Extragalactic Survey) and DEVILS (Deep Extragalactic Visible Legacy Survey). While HST was never intended to undertake large area surveys, through extensive quality control and filtering, we were able to extract a reliable and representative sample of fields distributed across the sky. Our final catalogs cover a combined , with individual filters covering areas ranging from . The combination of numerous independent sight-lines and area coverage allows us to reduce cosmic variance uncertainties in deep number counts to 0.21%-1.8%. For the first time we are able to establish a measurement of the IGL, , at 0.59 m using HST data. We obtain a cosmic optical background value of . Different techniques used to measure the COB, both directly and indirectly, have recently converged indicating that the COB arises almost exclusively from processes within galaxies. This in combination with the recent values reported from New Horizons and very high energy (VHE) constraints leaves very little room for any diffuse emission coming from outside the Milky Way.
keywords:
cosmology: cosmic background radiation, galaxies: statistics, methods: data analysis, cosmology: observations1 Introduction
@@
The extragalactic background light (EBL) is composed of all radiation originating from outside of our galaxy (Hill et al., 2018; Driver, 2021). The EBL is typically broken down into wavelength regimes such as the cosmic optical background (COB ) , cosmic infrared background (CIB ) and cosmic radio background (CRB); e.g. (Hill et al., 2018). The cosmic microwave background (CMB) is distinct from other astrophysical backgrounds as it is a relic from when the early Universe cooled to a temperature where neutral atoms could form. The remaining EBL from -ray to radio wavelengths contains information about all physical processes that have produced photons since recombination and thus can be considered separate from the CMB. The COB and CIB dominate the EBL and provide roughly equal energy contributions (Driver et al., 2016b; Koushan et al., 2021), though they both respectively fall a factor of 40x below the intensity of the CMB. The COB is perhaps the most important component of the EBL due to its intimate relationship with the cosmic star formation history (CSFH). The CSFH is defined as the volume averaged rate at which stars form as a function of either look-back time or redshift, for a thorough review see (Madau and Dickinson, 2014). The COB can be directly predicted from the CSFH assuming an implementation of a stellar initial mass function (IMF), dust attenuation model, metallicity evolution history, and stellar population model which has been demonstrated in works such as Andrews et al. (2017) and Koushan et al. (2021). Thus, empirical constraints of the COB can be used to test these models. The contribution from AGN is also important; especially at UV and mid-IR wavelengths, and can be modelled. The cosmic AGN luminosity history (CAGNH) is known to co-evolve with the CSFH (D’Silva et al., 2023). Despite its importance, the COB remains one of the most difficult portions of the EBL to study in large part due to the presence of foregrounds which outshine the EBL, see (Bernstein et al., 2002; Windhorst et al., 2022; Caddy et al., 2022; Carleton et al., 2022; Windhorst et al., 2023).
@@
Indirect detection techniques have become possible with advancements in the ability to study high energy -ray sources. The interaction between photons from distant -ray sources, typically blazars, and optical EBL photons results in an attenuation of TeV energy -rays through particle/anti-particle pair production interactions as discussed in Desai et al. (2017). Using blazar spectra with established redshifts, they are able to estimate the attenuation of -ray spectra by EBL photons along the line of sight between the source and detectors. While difficult, this method (hereafter VHE) is less susceptible to foregrounds and is also unaffected by issues such as cosmic variance and zero point uncertainties, which plague the IGL. However VHE techniques typically apply a universal spectral shape to the -ray sources and infer deviations from this spectrum as entirely due to interactions with EBL photons or intergalactic magnetic fields. The technique typically requires an existing COB model and just constraints the normalization of the model. Recent developments in VHE studies of the EBL, (e.g. Gréaux_2023; Gréaux_2024) have adopted a Bayesian framework to avoid limitations introduced by assuming a model and spectral shape for both the EBL and gamma ray source. Using the observed -ray spectra of blazars they present preliminary results in which they now extract the COB shape and amplitude. Uncertainties in measurements of the COB established by VHE studies to 15% and will be developed further in the coming years.
A commonly employed method to place constraints on the EBL without accounting for foregrounds is to use galaxy number counts and measure the total intensity of all resolved galaxies in works such as (e.g. Madau and Pozzetti, 2000; Totani et al., 2001; Xu et al., 2005; Dole et al., 2006; Hopwood et al., 2010; Keenan et al., 2010; Berta et al., 2011; Béthermin et al., 2012; Driver et al., 2016b; Koushan et al., 2021; Windhorst et al., 2022, 2023, Carter et al. 2025(submitted)). The IGL can be well constrained by surveying a large enough volume to obtain a representative sample of the Universe. Galaxy number counts, or rather the galaxy flux density, can be integrated to provide a measurement of their combined intensity. The IGL is insensitive to diffuse foregrounds as positive sky values across the image are subtracted during processing and/or source detection. In particular, for nearly 30 years, standard HST data processing has removed foreground sky levels by subtracting a 2D model fit to the image after initially detected discrete objects are masked out (Windhorst et al., 2022). The sky subtraction includes signals from the zodiacal light, diffuse galactic light, and scattered sun and/or moonlight. Similar processing is done for ground-based observations which must also deal with atmospheric effects such as airglow.
IGL based measurements of the COB are not sensitive to proposed diffuse sources of the EBL such as intra-cluster light (ICL), diffuse light from galaxy halos, or ultra-diffuse galaxies (UDGs) in galaxy clusters. Extended low surface brightness (SB) features may be lost during sky subtraction or remain undetected by automated source finding tools. In clusters, ICL may contribute as much as of cluster luminosity at SB levels fainter than (Montes, 2019). Similar fractions of from the extended halos of individual galaxies have been proposed in works such as (Ashcraft et al., 2018; Borlaff et al., 2019; Cheng et al., 2021). While missed by IGL studies, though as direct, indirect, and source-count based measurements improve tighter constraints will be imposed on these diffuse contributions to the COB.
@@
While establishing robust lower limits, the IGL is prone to uncertainties introduced by cosmic variance and magnitude zero-point differences present between the facilities providing the number counts. Cosmic variance and zero point errors remain the dominant sources of uncertainty in measuring the IGL. In works such as Driver et al. (2016b) these contribute uncertainties of 10% and 5%, respectively. More recent studies such as Koushan et al. (2021), through the use of larger area surveys and updated zero point calibrations, were able to bring these uncertainties down to 4%. However, Driver et al. (2016b) and Koushan et al. (2021) relied only on a small number of deep HST fields. Work such as Windhorst et al. (2011) and Rafelski et al. (2015) provide the faintest optical galaxy number counts and have only been surpassed by JWST recently in the NIR (Windhorst et al., 2023). The two HST surveys included in Driver et al. (2016b) and Koushan et al. (2021) cover a combined 40-60 arcmin2, and thus cosmic variance remains the dominant source of uncertainty in these surveys at 10-15% as estimated using the formula developed in Driver and Robotham (2010). Uncertainties in the IGL have also been reduced to between by Carter et al. 2025 (submitted). SKYSURF aims to use a significant fraction of the area covered by HST over 20 years of observations to further reduce uncertainties in measurements of the IGL primarily by reducing the cosmic variance errors of individual HST deep surveys used in previous works.
@@
Despite efforts to measure the COB both directly and through the IGL, disagreements between the upper limits established by work of Carleton et al. (2022) O’Brien et al. (2023), Mattila et al. (2017), Lauer et al. (2022) and lower limits established by the IGL in Driver et al. (2016b) and Koushan et al. (2021) remained factors of apart. The newest direct measurements of Postman et al. (2024) from New Horizons suggest that the tension between the IGL and direct measurements has finally been resolved. Again using the LORRI instrument aboard New Horizons, measurements have yielded a significant detection of the EBL at the LORRI pivot wavelength of , with the main change in Postman et al. (2024) being an improved method for estimating the DGL. Within their respective errors, the measurements of Postman et al. (2024), Koushan et al. (2021), and preliminary VHE measurements from Gréaux_2023 have finally converged. The explanation posed in Postman et al. (2024) is that the IGL composes the most significant component of the COB and the majority of extragalactic optical flux originates from within galaxies. The JWST work of Windhorst et al. (2023) suggests that no more than 20% of the COB can originate from outside of galaxies. As such, efforts to reduce uncertainties in the IGL can better constrain models and provide a near complete census of photon production from baryon-photon decoupling until the present day.
@@ 2.1 The WAVES Survey
WAVES is a spectroscopic redshift survey to be conducted using the European Southern Observatory’s 4-Metre-Multi-Object Spectroscopic Telescope (4MOST) (de Jong, 2011). In this work, bright galaxy number counts are produced in preparation for the WAVES survey (Driver et al., 2016a, 2019, Bellstedt et al. in prep) which covers a region of the sky. The galaxy catalogs used for WAVES target selection (Bellstedt et al. in prep) are produced from imaging taken from the VISTA KIlo-degree Infrared Galaxy (Edge et al., 2013, VIKING) and the VST Kilo-Degree Survey (Kuijken et al., 2019, VST-KiDS), both of which are ESO public data sets. The data was WCS aligned using ProPane (Robotham et al., 2023), and object detection is done using an rizY inverse variance stack using ProFound (Robotham et al., 2018). All bright galaxies were inspected for fragmentation following the detection phase and fixed. Following the detection phase, fluxes in each band were determined using forced aperture photometry from the detection stack. Stars brighter than GAIA g mag were masked. Star-galaxy separation was achieved using an unsupervised machine learning technique developed in Cook et al. (2024). WAVES galaxy number counts also include an object classification uncertainty. Objects are designated as either a star, galaxy, or ambiguous. Galaxy number counts were constructed both with and without the inclusion of ambiguous objects. The final galaxy counts are defined as the mean between the two and the uncertainty spans the range between the galaxy and galaxy + ambiguous counts. See Bellstedt et al. (in prep) for full details on these steps. In this work we use WAVES input catalog v1.2 with 7 bands (ugrizYJH) that overlap with the HST filter system. The completeness limit of each band is shown in Table 1 where completeness limits were defined by stepping back by one half-magnitude bin from where the counts reach a maximum.
@@ 2.2 The DEVILS Survey
For intermediate magnitude number counts we turn to DEVILS (Davies et al., 2021), using all three DEVILS fields, D02, D03, and D10. The three fields are centred on well-studied regions, XMM-LSS, ECDFS, and COSMOS, respectively. The DEVILS galaxy number counts are complete to AB 24 mag across all 3 fields and cover a combined area of 6 . A brief summary of areas and completeness limits are provided in Table 1. We elect to use the same 7 bands chosen from the WAVES survey, i.e. , . We note that D02 and D10 were observed by the Hyper Suprime Cam (HSC) (Aihara et al., 2017) in griz with the YJH bands observed by VISTA. In D03 the griz imaging was obtained by the VST. Finally, u-band imaging was obtained from the Canada-France-Hawaii-Telescope (CFHT) and VST.
@@ 2.3 The SKYSURF HST Archival Program
SKYSURF is an HST cycle 27-29 Legacy Archival program (AR-09955 & AR-15810) designed to use a significant amount of HST imaging data to measure the sky-surface brightness (sky-SB) across the wavelength range covered by HST, i.e., . This requires the reprocessing of archived HST data using consistent software pipelines throughout. During the last two years, the SKYSURF project has measured the sky-SB and divided it into its components, the ZL, DGL, and EBL, (e.g. O’Brien et al., 2023; Carleton et al., 2022; Windhorst et al., 2022, Carter et al. 2025 (Submitted)). The SKYSURF project is split into two separate teams using independent processing pipelines and methodologies to obtain mosaics, sky-SB values, and object catalogs. This paper constitutes the portion of the project led from the University of Western Australia focused on integrated galaxy light (IGL) measurements using object catalogs generated using the ProPane and ProFound software packages, (Robotham et al., 2023, 2018) as was used by WAVES and DEVILS. In Carter et al. 2025 (submitted), an independent processing of the SKYSURF archive was performed using the AstroDrizzle (Gonzaga et al., 2012) for image alignment and stacking. Source detection was done using SourceExtractor (Bertin and Arnouts, 1996). Our overriding objective is to produce reliable galaxy number counts across as broad a magnitude and wavelength range as possible. We begin our SKYSURF selection by only including HST imaging with an exposure time of 200 seconds or longer and at galactic latitude . A projection showing the location of SKYSURF data used in this work is shown in Figure 1. All data was downloaded from the Mikulski Archive for Space Telescopes (MAST)111https://mast.stsci.edu/portal/Mashup/Clients/Mast/Portal.html.
A number of tests and findings from earlier SKYSURF papers have contributed to the final decisions and methodology used in this work. Our science goals require an ability to demonstrate reliable sky surface brightness estimation and subtraction. In SKYSURF I Windhorst et al. (2022), we investigated nine different sky background estimation algorithms. We produced simulated WFC3IR images with known sky values either with or without an input gradient across the image. A number of different observing conditions such as sun angle, moon angle, or off-axis bright stars can introduce a non-uniform sky value which varies across the image.
We found that the ProFound based sky estimation was able to recover input sky values within & % for images with and without sky gradients, respectively. SKYSURF IV (O’Brien et al., 2023) implemented different sky estimation techniques on SKYSURF data, demonstrating that sky background levels increase as ecliptic latitude and/or sun angle decrease. This is not a surprising result, but reinforces the need to reliably subtract the sky background dominated by zodiacal light.
@@ 3.1 Galactic Extinction Corrections
We correct for galactic extinction using the Planck satellite dust maps accessible through the web tool333https://irsa.ipac.caltech.edu/applications/DUST/ with extinction estimates provided by Finkbeiner et al. (1998). Given the small angular size of SKYSURF fields compared to the resolution of the dust maps, we elect to use the centre of the field to obtain a single E(B-V) value and thus apply one extinction value to all objects in that field. Using the Spanish Virtual Observatory tool provided by (Rodrigo and Solano, 2020)444http://svo2.cab.inta-csic.es/theory/fps/, we obtain values for each HST filter. The values adopted for each HST filter used in this work along with the pivot wavelengths adopted in this work are listed in Table 3. is defined as the ratio between extinction at . and visual extinction. In Figure 6 we show the Galactic extinction values for each mosaic as a function of Galactic latitude. By design, SKYSURF omitted all observations within 10 degrees of the galactic plane, as indicated by the vertical lines in Figure 6.